What is the period of the orbit? 13 Telmoved, what is the radius of the electron or 23, I Singly ionized (one electron removed) atoms are acce then passed through a velocity selector consisting of perpendia electric and magnetic fields. The electric field is 155 V/mai magnetic field is 0.0315 T. The ions next enter a uniform mag pendicular and the n magnetic
feld of magnitude 0.0175 T that is oriented perpendicular to their velocity. (a) How fast are the ions moving when they emerge from the velocity selector? (b) If the radius of the path of the ions in the second magnetic field is 17.5 cm, what is their mass? Ancient meat eating. The amount of meat in prehistoric diets can BIO be determined by measuring the ratio of the isotopes nitrogen-15 to nitrogen-14 in bone from human remains. Carnivores concentrate 15N, so this ratio tells archaeologists how much meat was consumed by ancient people. Suppose you use a velocity selector to obtain singly ionized (missing one electron) nitrogen atoms of speed 8.50 km/s and bend them along a semicircle within a uniform mag- netic field. If the 14N atoms travel along a semicircle with a diameter of 29.2 cm, find the separation of the 1'N and 15N isotopes at the de- tector. The measured masses of these isotopes are 2.32 x 10-26 kg (I4N) and 2.49 x 10 26 kg (15N).
Answer #1

a] The velocity selector is such that the electric and magnetic forces on the ion are balanced.

qE = qvBsin90

=> E = vB

=> v = E/B = 4920.63 m/s

this is the speed of the ions when they emerge from the velocity selector.

b] The force due to the magnetic field = centripetal force required for circular motion

mut quB =

=> r = \frac{mv}{qB}

here, r = 0.175 m

so, m = 9.958 x 10-26 kg = 59.988 amu

this is the mass of the ion.